### v1

@@ 3 Implementation
In this section, we show the simulation results of several different configurations; as described in section 2.1. This numerical study is designed to simulate the experiments conducted by Glas and Westwater (1964). The experimental result confirm the driven force of bubble growth is diffusion and measure the bubble growth rate. The comparsion between following numerical and experimental growth rate will verify the accuracy of simulation. The physical properties corresponding to an alkaline solution (typically used in industrial water electrolysis) are shown in Table 2.
The simulation results are presented with dimensional units for comparison with experimental measurements. According to the analytical and experimental work, the growth of electrochemically generated bubbles follows the same functional relationship as the solution for bubble growth in a supersaturated liquid described by the Scriven model. Different stages or regimes for bubble growth have been characterized by a power law: growth controlled by inertia ($R\propto t$), by diffusion ($R\propto t^{1/2}$) or by reaction limitation ($R\propto t^{1/3}$) , see (Angulo et al., 2020). The growth behavior ($R\propto t^{1/2}$) reflects a standard analytical solution for diffusive bubble growth. To validate the specific relationship, we perform a series of axisymmetric simulations first.
@@ 3.1 Axisymmetric simulation validation
In this section, we assume there is only one bubble (with an initial diameter $D_{b}$) on the cathode. In fig. 5, the left-hand side of the snapshots illustrates hydrogen evolution during the bubble’s growth stage, where color represents the dissolved hydrogen concentration in the liquid phase. For visualization consistency, the gas phase inside the bubble will also be colored, but will not be calculated (based on the assumption of constant pressure inside the bubble) and will remain at 1. The same map is applied to the rest of the snapshots in the following. Hydrogen is produced at the electrode surface, so hydrogen saturation near the electrode wall is more significant than that in the bulk liquid. Completing the visualization, the right side of the snapshots displays the symmetric grid cell and interface, clearly demonstrating the mesh refinement near the gas-liquid interface.
The growth rate of a single bubble confirms the result of the diffusion-controlled power law ($R\propto t^{1/2}$). As shown in fig. 6, a mesh independence study demonstrates that mesh refinement level 8 will be enough for interface tracking. A logarithmic plot of the bubble growth for different current densities between $t=0.02\mathrm{s}$ and $t=0.2\mathrm{s}$ of our simulation (mesh refinement level 8) is shown in fig. 7.
In our simulation case, the hydrogen flux is applied before the bubble nucleation time ($t_{n}=20\mathrm{ms}$), with the average hydrogen concentration on the electrode wall increasing in time. After nucleation, the concentration difference drives bubble growth through the diffusion process. As a result, the bubble growth behavior shows some discrepancies from the analytical solution during the early stages of diffusion. For smaller current density $I=100\mathrm{A/m^{2}}$, the growth exponent is smaller than $1/2$. For much larger current density $I=1000\mathrm{A/m^{2}}$, the growth follows the power law ($R\propto t^{0.8}$) better. In a similar simulation using the immersed boundary method, see Khalighi et al. (2023), the growth exponent 0.8 is also observed in single bubble growth.
@@
Scriven’s result is valid for a spherically symmetric concentration field, bubble radius much larger than the initial radius, constant solubility, and, far away from the bubble, constant concentration and zero velocity. This solution is confirmed in cases where the thickness $\delta_{\mathrm{H_{2}}}$ of the diffusion boundary layer, surrounding the bubble is small compared with the diameter of the bubbles.
Given the constraints of the analytical solution, the discrepancies in our simulation can be explained. The hydrogen flux at the boundary determined by current density influences the growth exponent increasing the averaged hydrogen concentration over the electrode wall. At the beginning stage, the growth exponent is not close to $1/2$. The bubble radius is relatively small compared to the thickness of the diffusion boundary layer. For high current densities $I=1000\mathrm{A/m^{2}}$, the enhanced hydrogen flux elevates the supersaturation level and creates a strong diffusion boundary layer. The diffusion layer continuously provides $\mathrm{H_{2}}$ for bubble growth, which by consequence leads a much faster $\mathrm{H_{2}}$ production than consumption and a growth exponent exceeding $1/2$. Conversely, lower current densities reduce hydrogen availability, can not fulfill the consumption of bubble growth, leading to sub $1/2$ exponents.
@@
After the initial growth stage, as the bubble radius increases, the $\delta_{\mathrm{H_{2}}}$ is much smaller compared with the bubble radius. The growth exponents for different boundary condition are all approaching $1/2$, which indicates that the overall growth is diffusion controlled.
@@
Figure 5: The dissolved hydrogen concentration for a bubble growing at different times (The current density is $I=\mathrm{1000A/m^{2}}$).The color lengend represents the numerical $\mathrm{H_{2}}$ concentration.
Figure 6: Bubble radius growth for bubbles with $\theta=90\degree$ for different current densities. The units of $I$ are $\mathrm{A/m^{2}}$
Figure 7: The growth bubble radius slope and growth exponent variation.
To further compare the numerical results with experimental data, additional simulations are performed with a smaller contact angle ($\theta=35^{\circ}$), which is the same value from the experimental study(see Glas and Westwater (1964)). From the analytical solution $R=2\beta t^{1/2}$, the growth rate $\beta$ could be deduced. The results in terms of the influence of current density on the growth rate are reported in fig. 8. The first comparison indicates that the growth rate increases with rising current density, attributed to the high concentration of dissolved hydrogen near the electrode. This observation is supported by both the numerical simulations conducted by Gennari et al. (2022) and experimental results from Glas and Westwater (1964). The observation can be explained by the current density controlling the hydrogen flux at the electrode, which drives bubble growth.
Figure 8: Comparison of growth rate between experimental and numerical results.
We find that a bubble with a contact angle of $\theta=90^{\circ}$ grows more rapidly than one with a contact angle of $\theta=35^{\circ}$. This difference can be explained by considering the geometric difference of different contact angles. The mass transfer through diffusion occurs only through the interface between the gas and liquid phase. In the $\theta=90^{\circ}$ scenario, the bubble elongates more in the r-direction (parallel to the electrode wall), thereby increasing its exposure to the high-concentration region close to the electrode. So, the mass transfer for the bubble with $\theta=90\degree$ would be higher. A qualitative comparison between $\theta=90\degree$ and $\theta=35\degree$ is shown in fig. 9 and fig. 10.
It is notable that the simulation of a single bubble overpredicts the growth rate compared to the experimental measurements. This discrepancy is observed not only in the present simulation but also in the results from Gennari et al. (2022). In fact, there are at least four active nucleation sites on the electrode surface from the experiment of Glas and Westwater (1964). This suggests that experimental observations of single bubble growth are generally influenced by neighboring bubbles. Depending on the material and surface properties of the electrode, if multiple nucleation sites are active, the dissolved hydrogen in the liquid is consumed simultaneously by all of them. Consequently, the relative growth rates are smaller than in the case of a single nucleation site, where the entire amount of $\mathrm{H_{2}}$ produced by the electrode is available for a single bubble on the wall. This mechanism explains the overprediction in the simulations. To further investigate the effect of variations in nucleation site distribution, a 3D simulation would be necessary, as the scenario would no longer be axisymmetric.
Figure 9: Bubble shape and contour of $\mathrm{H_{2}}$ concentration for $\theta=90\degree$.
Figure 10: Bubble shape and contour of $\mathrm{H_{2}}$ concentration for $\theta=35\degree$.
3.2 3D simulation extension
@@
To extend the numerical scheme to 3D simulation, we set the domain as a cubic (fig. 3). And the boundary condition would be modified to suit the domain. The electrode will be placed in the middle of the domain bottom(XY plane), which will give the same boundary condition of hydrogen flux as axisymmetric simulation, see eq. 4. The top plane(XY) will be set as an outflow boundary condition, and the surrounding area (XZ and YZ planes) adopt a no-slip wall boundary condition.
@@
In this section, a single bubble is modeled in 3D to validate the 3D applicability of the numerical method. We compare the bubble growth prediction of the 3D simulation to that of the corresponding axisymmetric simulation. Simulations of configurations No.1 and No.7 (Table 1) with three different current densities are listed for comparison. The contact angle is $90\degree$. Same bubble growth slope from section 3.1 are observed in fig. 11. Since the convergence has been verified in section 3.1, the good agreement with the axisymmetric simulation results proves the compatibility of 3D simulation.
Figure 11: Bubble radius growth for bubbles with $\theta=90\degree$ for different current densities. The units of $I$ are $\mathrm{A/m^{2}}$.
@@ 4.1 Multiple bubbles growth
There is a notable decrease in the growth rate and exponent of hydrogen bubbles when multiple nucleation sites are active from the experiment of Glas and Westwater (1964). To reproduce this key observation, this section numerically investigates the mutual bubble interactions for multiple nucleation sites. Instead of a single bubble, we model 2 bubbles and 4 bubbles growing at the electrode. To highlight the effect of nucleation sites, a higher boundary flux is applied, with the current density set to $\mathrm{1000A/m^{2}}$. The simulation results for the average bubble radius growth of two and four bubbles are presented in fig. 12. The growth of multiple nucleation sites is mutually suppressed in the stage before bubble coalescence. During this stage, hydrogen is consumed by all the bubbles simultaneously, leading to a reduced growth rate and exponent. The more active nucleation sites, the slower the growth, as shown in the comparison between experimental and simulation results in fig. 13. However, an unexpected behavior of the growth seems puzzling. Intuitively, merged bubbles should approximate the growth dynamics of a single nucleation site, leading to convergent growth across all cases. Contrary to this assumption, even after bubble coalescence($t>t_{m}$), differences in the growth persist.
Figure 12: Bubble growth of configurations No.7 to No.9. The current density is $I=\mathrm{1000A/m^{2}}$ and the slope represents the growth exponent.
Figure 13: Influence of nucleation sites on the growth rate.
A possible explanation lies in the relative position of bubbles within the hydrogen concentration layer. As shown in fig. 12, the trend remains consistent: the more active nucleation sites, the slower the growth, regardless of the stage (before or after bubble coalescence).
@@
Figure 14: Comparison of bubble growth between different nucleation sites. The current density is 1000 A/m2. The view direction is the same as in fig. 3. The hydrogen contour on the intersecting plane (YZ) in the middle of the bubble is transferred to the back plane.
The key reason for the different growth behavior after coalescence is the bubble radius. Observations indicate that the radius of the merged bubble in the four-nucleation-site case is larger than in the dual-site case, while the bubble in the single-site case has the smallest radius. Despite this, the hydrogen concentration layer remains nearly identical across different nucleation-site cases at the same time. This is because the hydrogen flux at the boundary is sufficiently high, making the amount of hydrogen consumed by bubble growth negligible. Then, the crucial factor is immersion depth: bigger bubbles immersed shallower into the hydrogen concentration layer, leading to a lower mass transfer rate, which means a smaller growth exponent, as illustrated in fig. 12.
Supporting evidence emerges from direct comparisons: at the same time $t_{m}$, the four-nucleation-site configuration produces marginally larger bubbles than the single-site case, see fig. 14. However, the constant hydrogen flux induces a countervailing effect as illustrated in the snapshots fig. 14, the bubble surface is colored by the mass transfer rate. A higher mass transfer rate slows the increase of the hydrogen concentration boundary layer for the four-nucleation-site case, enabling shallower bubble immersion after merge. This feedback mechanism drives a gradual decrease in growth exponent, ultimately causing convergence across nucleation-site configurations. As shown in fig. 12, the growth exponent gradually approaches a slope of 0.5.
@@
To quantitatively assess mass transfer, we analyze the time-dependent bubble Sherwood number defined in eq. 8, which quantifies mass transport efficiency to the bubble. The initial condition $\zeta=1$, delay the plot until $t>t_{n}$, which is shown in fig. 15. It should be noticed that multiple nucleation sites exhibit elevated $\mathrm{Sh_{b}}$ initially. A decrease occurs after bubble merge($t_{m}$), which causes the $\mathrm{Sh_{b}}$ to be smaller than the one for the single nucleation site. The plot of $\mathrm{Sh_{b}}$ cross-verified the explanation of growth exponent above. First, the mutual suppression. While total mass transport increases with nucleation sites, individual bubble transport efficiency is inhibited through competitive $\mathrm{H_{2}}$ consumption. Second, the immersion depth: lower after-merge $\mathrm{Sh_{b}}$ directly correlates with diminished mass transfer rates for multiple nucleation sites, consistent with the immersion depth arguments.
Figure 15: Corresponding Sherwood number for simulations of configurations No.7 to No.9.

### v4

@@ 3 Implementation
In this section, we show the simulation results for several different configurations, as described in Section˜2.1. This numerical study is designed to simulate the experiments conducted by Glas1964. The experimental results confirm that the driving force for bubble growth is diffusion and measure the bubble growth rate. The physical properties corresponding to an alkaline solution (typically used in industrial water electrolysis) are shown in Table˜2.
It should be noted that the original surface tension $\sigma=0.075\mathrm{N/m}$ is decreased by a factor of $10^{-4}$ and the hydrogen density $\rho_{d}=0.08\,\text{kg/m}^{3}$ is increased to $\rho_{d}=0.8\,\text{kg/m}^{3}$. The molar mass is scaled by the same factor, ensuring that the relative volume change remains unaffected ($\Delta V\propto M/\rho_{d}$). Sensitivity tests are therefore required to assess the effects of the decreased surface tension and increased hydrogen density.
Decreasing the surface tension is introduced to reduce computational cost, as the surface-tension scheme imposes a restrictive time-step constraint. A comparison of simulations with modified $\sigma$ in Fig.˜4 indicates that increasing surface tension has only a minor influence on bubble growth. An analysis of the Weber number for different values of $\sigma$ further shows that surface tension remains the dominant force during the bubble growth stage, even after $\sigma$ is decreased. Bubble detachment, however, is significantly more sensitive to the choice of $\sigma$, as reflected by the Bond number. To maintain consistency with the theoretical prediction of the detachment radius, the same decreased value of $\sigma$ is adopted.
Regarding hydrogen density, its physical value is relatively small, leading to a large density ratio $(\rho_{c}/\rho_{d}=12450)$, which would be a problem for numerical simulations. A high density ratio is observed to slow the convergence of the multigrid solver. Sensitivity tests for different hydrogen densities, shown in Fig.˜4, demonstrate that increasing $\rho_{d}$ has a negligible effect on bubble growth. Based on the results above, we can confirm that a decrease of surface tension $\sigma$ by a factor of $10^{-4}$ and an increase of hydrogen density to $10\times\rho_{d}$ provide an acceptable balance between numerical stability, computational efficiency and physical accuracy.
Figure 4: Sensitivity test on decreased surface tension (a), and increased hydrogen density (b).
The simulation results in this work are presented in dimensional units to facilitate comparison with experimental measurements. According to the analytical and experimental work, the growth of electrochemically generated bubbles follows the same functional relationship as the solution for bubble growth in a supersaturated liquid described by the Scriven model. Different stages or regimes for bubble growth have been characterized by a power law: growth controlled by inertia ($R\propto t$), by diffusion ($R\propto t^{1/2}$) or by reaction limitation ($R\propto t^{1/3}$) (Angulo2020). The growth behavior ($R\propto t^{1/2}$) reflects a standard analytical solution for diffusive bubble growth. To validate the specific relationship, we first perform a series of axisymmetric simulations on bubble growing at the contact angle of $\theta=90\degree$.
@@ 3.1 Axisymmetric simulation validation
In this Section, we assume there is a single bubble (with an initial diameter $D_{b}$) on the cathode. In Fig.˜5, the left-hand side of the snapshots illustrates hydrogen evolution during the bubble’s growth stage, where color represents the dissolved hydrogen concentration in the liquid phase. For visualization consistency, the gas phase inside the bubble will also be colored, but will not be calculated (based on the assumption of constant pressure inside the bubble) and will remain at $1\mathrm{mol/m^{3}}$. The same color map is applied to the rest of the snapshots in the following. Hydrogen is produced from the electrode surface, and the saturation near the electrode wall is more significant than that in the bulk liquid. Completing the visualization, the right side of the snapshots displays the grid and the moving interface, clearly showing adaptive quadtree mesh refinement.
A mesh independence study has been performed to demonstrate that a mesh refinement level of 8 will be sufficient for accurate resolution. The following logarithmic plot Fig.˜6 presents bubble growth for five different current densities between $t_{n}=0.02\,\mathrm{s}$ and $t=0.2\,\mathrm{s}$. In our simulation case, the hydrogen flux is applied before the bubble nucleation time ($t_{n}=20\mathrm{ms}$), with the average hydrogen concentration on the electrode wall increasing in time. After nucleation, the concentration difference drives bubble growth through the diffusion process.
The inertia-driven and diffusion-driven growth regimes correspond to slopes of 1 and $1/2$, respectively. As shown in Fig.˜6, the transition to diffusion-driven growth is evident, but the expected inertia-driven regime is missing. The steepest slope is observed only during the initial growth stage at the highest current density, and is close to $0.8$.
@@
The diffusion-driven growth slope of $1/2$ is assumed to be valid for a spherically symmetric concentration field, a bubble radius much larger than the initial radius, constant solubility, and, far from the bubble, constant concentration and zero velocity. This solution is confirmed in cases where the thickness $\delta_{\mathrm{H_{2}}}$ of the diffusion boundary layer surrounding the bubble is small compared with the diameter of the bubbles. Given these constraints of the analytical solution, the discrepancies in our simulation can be explained. The hydrogen flux at the boundary, which is determined by current density, influences the growth exponent by increasing the averaged hydrogen concentration on the electrode wall. At the beginning stage, the growth exponent is larger than $1/2$. The bubble radius is relatively small compared to the thickness of the diffusion boundary layer.
At high current densities, such as $I=1000\,\mathrm{A/m^{2}}$, the increased hydrogen flux raises the supersaturation level and generates a strong diffusion boundary layer. This layer continuously supplies $\mathrm{H_{2}}$ for bubble growth, resulting in a production rate that surpasses consumption and yields a growth exponent greater than $1/2$. Numerically, the stronger flux produces a steeper $\mathrm{H_{2}}$ concentration gradient, which enhances mass transfer across the interface. In contrast, lower current densities limit hydrogen availability, making it insufficient to sustain bubble growth and leading to sub-$1/2$ exponents. The decreasing trend of the growth exponent with lower current densities confirms this behavior.
@@
After the initial growth stage, as the bubble radius increases, the $\delta_{\mathrm{H_{2}}}$ is much smaller compared with the bubble radius. The growth exponents for different boundary conditions are all approaching $1/2$, which indicates that the overall growth is diffusion-controlled.
@@
Figure 5: Frames show the dissolved hydrogen concentration for a bubble growing at the contact angle of $\theta=90\degree$(with the current density of $I=\mathrm{1000A/m^{2}}$). The color legend represents $\mathrm{H_{2}}$ concentration, while the upper limit is $0.2\mathrm{mol/m^{3}}$.
Figure 6: The bubble radius growth slope and growth exponent variation.
3.2 Three-dimensional simulation
@@
To extend the numerical scheme to three-dimensional simulation, we set the domain as a cube (Fig.˜3). And the boundary condition would be modified to suit the domain. The electrode will be placed in the middle of the domain bottom boundary ($xy$ plane), which will give the same boundary condition of hydrogen flux as an axisymmetric simulation, see Eq.˜4. The top $xy$ plane will be set as an outflow boundary condition, and on the lateral boundaries ($xz$ and $yz$ planes), a no-slip boundary condition is set.
@@
A single bubble is modeled in three dimensions to validate the applicability of the numerical method. We compare the bubble growth prediction of the three-dimensional simulation to that of the corresponding axisymmetric simulation. Specifically, simulations of configurations No. 1 and No. 6 (Table˜1) are compared. The results show that the bubble growth slopes obtained from the axisymmetric simulations are identical to those from the three-dimensional simulation. Since the numerical convergence has been verified in Section˜3.1, this good agreement confirms the consistency of the three-dimensional simulation. Consequently, the three-dimensional configuration is used for the subsequent simulations.
@@ 4 Result and discussion
4.1 Single bubble growth
Based on the experiments of Glas1964, both the current density and the electrode surface covering (determined by the contact angle $\theta$) influence the bubble growth. A notable observation is that the contact angle is variable. Glas1964 found that the change is large, from about 70°to 20°. However, for simplicity, our numerical model assumes a constant contact angle throughout the bubble growth evolution. For further comparison, another series of simulations is performed with a smaller contact angle ($\theta=35^{\circ}$), which corresponds to a hydrophilic electrode surface.
From the analytical solution $R=2\beta t^{1/2}$, the growth rate $\beta$ can be determined. The fitting method used by Glas1964 involves plotting the bubble radius $R$ versus $t^{1/2}$, from which $\beta$ is computed directly from the slope. Using the same historical fitting approach, the numerical results showing the influence of current density on the growth rate are presented in Fig.˜7. Because the bubble growth exponent does not perfectly match $1/2$ during the early growth stage (as discussed in Fig.˜6), we performed three local fittings at different times for each case, with the resulting error bar indicating the variability among these estimates. The results clearly show that the growth rate increases with increasing current density. This trend is expected, as the current density directly dictates the hydrogen flux at the electrode surface, which drives bubble growth through higher concentrations of dissolved hydrogen near the electrode.
Figure 7: Comparison of growth rate of a single bubble between experimental and numerical results.
Then, we find that a bubble with a contact angle of $\theta=90^{\circ}$ grows faster than one with a contact angle of $\theta=35^{\circ}$. This difference can be explained by considering the influence of the contact angle on the bubble shape and placement. Mass transfer through diffusion occurs only through the interface between the gas and liquid phases. In the $\theta=90^{\circ}$ scenario, the bubble elongates more along the electrode wall, thereby increasing its exposure to the high-concentration region close to the electrode. As a result, the effective diffusive mass transfer to the bubble with $\theta=90\degree$ is enhanced.
Furthermore, Fig.˜8 presents the growth rate trend deduced from several different contact angles. The simulations cover the range from $35^{\circ}\leq\theta\leq 90^{\circ}$ for the same initial nucleation diameter. This range is relevant as most electrode surfaces are designed to be hydrophilic to facilitate the removal of attached bubbles. The results indicate that the bubble-growth coefficient generally increases with the contact angle, until the upper limit of our simulations of $\theta=90^{\circ}$. However, as the contact angle approaches the lower limit, the bubble-growth coefficient is no longer dominated by the contact angle. In some instances, such as at a current density of $I=500\,\mathrm{A/m^{2}}$, this trend is even reversed. While we have no explanation for this reversal, we attribute the weakening influence of the contact angle near the lower limit to a geometrical effect. Indeed, as the contact angle decreases, the bubble geometry changes only a little while it approaches a perfect sphere, and the influence of the contact angle on growth becomes less significant. Consequently, simulation results for contact angles smaller than $\theta=35^{\circ}$ are not shown here.
Figure 8: Comparison of growth rate of a single bubble between experimental and numerical results for various contact angles.
A significant discrepancy in the effect of current density is also noted when comparing the simulated and experimentally measured bubble-growth coefficients $\beta$, as shown in Fig.˜7. For the $\theta=35^{\circ}$ case, the numerical model underestimates $\beta$ at low current densities. The deviation is more complex for $\theta=90^{\circ}$, where $\beta$ is underestimated at low current densities but is conversely overestimated at higher current densities. This discrepancy is observed not only in the present simulation but also in the results of Gennari2022.
The underestimation is likely due to the difference between the numerical setting and the experimental conditions. Glas1964 established a nonlinear relationship between the growth coefficient $\beta$ and the current density $I$, as the driving force (local supersaturation) inherently depends on the current density. This model $\beta=I^{a},a<1$ effectively describes the experimental minor change of $\beta$ observed for $(100<I<500\,\mathrm{A/m^{2}})$. In contrast, our numerical work assumes that the rate of $\mathrm{H_{2}}$ production is directly proportional to the electrical current. This simplified setting cannot perfectly replicate the real situation, where bubble production at any site is intermittent in practice. Furthermore, the simulation uses a typical nucleation time of $0.02\mathrm{s}$. This time point determines the concentration layer of dissolved hydrogen when bubble growth begins. However, the nucleation time in experiments can vary between $0.02\mathrm{s}$ to $0.1\mathrm{s}$ (Glas1964). The difference in nucleation time at low current densities results in a weaker hydrogen concentration boundary layer than the experimental boundary layer, and therefore a smaller growth coefficient.
To explain the overestimation observed in the simulations, two major differences between the simulated and experimental conditions should be noted. First, the contact angle is found to change from $70\degree$ to $20\degree$ in the experiment. Since our simulations utilize a constant contact angle, a perfect match across all experimental conditions is not expected. However, it is noteworthy that the discrepancy is minimal when comparing the simulation results within the experimentally observed range of $35\degree\leq\theta\leq 70\degree$.
Second, obtaining a truly isolated, single active nucleation site on an electrode surface is challenging in both modern experiments and earlier studies (Glas1964). This suggests that experimental measurements of presumed single-bubble growth may, in fact, be influenced by neighboring bubbles, such as a bubble carpet or multiple nucleation sites. When multiple nucleation sites are active (a condition dependent on the electrode material and its surface properties), the dissolved hydrogen in the liquid is consumed simultaneously by several bubbles. Consequently, the individual growth rates are reduced compared with the idealized case of a single nucleation site, which benefits from the full hydrogen flux from the electrode, contributing to the growth of one bubble alone.
@@ 4.2 Multiple bubbles growth
There is a notable decrease in the growth rate and exponent of hydrogen bubbles when multiple nucleation sites are active (Glas1964). This Section numerically investigates the mutual bubble interactions for multiple nucleation sites. Instead of a single bubble, we model two bubbles and four bubbles growing at the same electrode. Since the influence of the contact angle on bubble growth has already been discussed in Section˜4.1, the results presented in this section focus on bubbles growing at a contact angle of $\theta=90\degree$.
The growth of multiple nucleation sites is mutually suppressed in the stage before bubble coalescence from the growth slope Fig.˜9. During this stage, hydrogen is consumed by all the bubbles simultaneously, leading to a reduced growth rate and exponent. The larger the number of active nucleation sites, the slower the growth. However, an unexpected behavior of the bubbles’ growth seems puzzling. Intuitively, once the bubbles merge, they should approximate the growth dynamics of a single bubble case. Contrary to this assumption, even after bubble coalescence($t>t_{m}$), differences in the growth persist for a period. Then the growth curves of all cases converge, as shown in Fig.˜9.
A possible explanation lies in the relative position of bubbles within the hydrogen concentration layer. As shown in Fig.˜9, the trend remains consistent: the more active nucleation sites, the slower the growth, regardless of the stage (before or after bubble coalescence).
Figure 9: Growth of multi-bubbles at the contact angle of $\theta=90\degree$. The current density is $\mathrm{1000A/m^{2}}$ and the slope represents the growth exponent.
The key reason for the different growth behavior after coalescence is the bubble radius. Observations indicate that the radius of the merged bubble in the four-nucleation-site case is larger than in the dual-site case, while the bubble in the single-site case has the smallest radius. Despite this, the hydrogen concentration layer remains nearly identical across different nucleation-site cases at the same time. This is because the hydrogen flux at the boundary is sufficiently high, making the amount of hydrogen consumed by bubble growth negligible. Then, the crucial factor is immersion depth: bigger bubbles immersed less deeply into the hydrogen concentration layer, leading to a lower mass transfer rate, which means a smaller growth exponent, as illustrated in Fig.˜9.
Supporting evidence emerges from direct comparisons: at the same time $t_{m}$, the four-nucleation-site configuration produces marginally larger bubbles than the single-site case, as shown in Fig.˜10.
However, the constant hydrogen flux induces a countervailing effect. As illustrated by the snapshots in Fig.˜10, where the bubble surface is colored by the local mass transfer rate, a higher overall mass transfer rate slows the increase of the hydrogen concentration boundary layer for the four-nucleation-site case. This effect enables shallower bubble immersion after coalescence. This feedback mechanism drives a gradual decrease in the growth exponent, ultimately causing convergence across different nucleation-site configurations. As shown in Fig.˜9, the growth exponent gradually approaches a slope of $1/2$.
@@
Figure 10: Frames show the comparison between the single bubble and four bubbles at the contact angle of $\theta=90\degree$. The current density is 1000 A/m2. The view direction is the same as in Fig.˜3. The hydrogen contour on the intersecting plane (yz) passing through the middle of the bubble is projected to the back plane. The bubble interface is colored according to the local mass-transfer rate.
@@
To quantitatively assess mass transfer, Fig.˜11 allows to analyze the time-dependent bubble Sherwood number defined in Eq.˜8, which quantifies mass transport efficiency to the bubble. It should be noticed that multiple nucleation sites exhibit elevated $\mathrm{Sh_{b}}$ initially. A decrease occurs after bubble coalescence($t_{m}$), which causes the $\mathrm{Sh_{b}}$ to be smaller than the one for the single nucleation site. The plot of $\mathrm{Sh_{b}}$ cross-verified the explanation of the difference in growth rate above. First, the mutual suppression. While total mass transport increases with nucleation sites, individual bubble transport efficiency is inhibited through competitive $\mathrm{H_{2}}$ consumption. Second, the immersion depth: lower after-merge $\mathrm{Sh_{b}}$ directly correlates with diminished mass transfer rates for multiple nucleation sites, consistent with the immersion depth arguments.
Figure 11: The Sherwood number for simulations with different numbers of nucleation sites. The current density is $1000\mathrm{A/m^{2}}$.
As shown in the comparison between experimental and simulation results in Fig.˜12, the growth rates obtained from the multiple-bubble cases exhibit better agreement with experimental data than those from the single-bubble case (Fig.˜7). This observation supports the earlier conjecture regarding the influence of neighboring bubbles. Although the reported experimental growth rate used for comparison was acquired for a single bubble attached to the electrode, likely, small bubbles forming underneath were not captured during measurement. Consequently, the growth rates predicted by the multiple-bubble simulations may more accurately represent the actual experimental conditions.
Furthermore, the experimental data itself exhibits a considerable standard deviation, and the growth coefficient $\beta$ was obtained through a fitting process without associated error bars. This experimental uncertainty presents an alternative explanation for the observed differences between simulation and experimental measurement.
Figure 12: Comparison of growth rate for multiple nucleation sites with the contact angle of $\theta=90\degree$.
Figure 13: Comparison of growth rate for single site and multiple nucleation sites with the contact angle of $\theta=90\degree$.
To comprehensively assess the influence of nucleation sites, Fig.˜13 compares the average bubble growth coefficients $\beta$ for one, two, and four bubbles. In addition to the expected suppression of the growth coefficient, the results reveal that the relative impact of multiple nucleation sites becomes more significant at higher current densities. This observation may help explain the earlier discrepancies between simulations and experiments: the higher current density amplifies the effect of other parameters, such as contact angle and the density of nucleation sites. Notably, the overestimation of the bubble-growth coefficient is remarkable at the current density of $I=1000\,\mathrm{A/m^{2}}$, whereas at lower current densities $(300<I<800\,\mathrm{A/m^{2}})$, the numerical predictions are comparable to the experimental measurements.